Daniel wants to rent a boat. The rental shop charges an initial fee of $15 plus $12 per hour to rent the boat.
A. If the total cost to rent the boat was $99, for how many hours was the boat rented?
B. If he rented the boat for 6 hours, how much would that cost?

Answer by mananth(16946)


f(c) = 15 + 12 (h)
99= 15+12h
84 =12h
h=7 Number of hours 7
f(c) = 15+(12)(6)
=87
$87 is the charge